OpenStudy (anonymous):

how do we type problems that have the square root symbol in it how do we type problems that have the square root symbol in it @Mathematics

OpenStudy (anonymous):

sqrt(whatever)

OpenStudy (unklerhaukus):

press option v

OpenStudy (anonymous):

option v??

OpenStudy (anonymous):

click on equation then u will see the options there only

OpenStudy (anonymous):

\[4\sqrt{256}\]

OpenStudy (anonymous):

64

OpenStudy (anonymous):

lol i can do it here but it wont copy and paste to the room

OpenStudy (anonymous):

its gonna take me forever to check my problems lol

OpenStudy (anonymous):

\[\sqrt[4]{}256\]

OpenStudy (anonymous):

here u got it right

OpenStudy (anonymous):

\[\sqrt[4]{?}256\]

OpenStudy (anonymous):

lol this is gonna be fun....

OpenStudy (anonymous):

can go to equation editor

OpenStudy (anonymous):

\[\sqrt[4]{256}\]

OpenStudy (anonymous):

i tried that from my word on my computer and it wont copy and paste to here either

OpenStudy (anonymous):

yea thats what it is sup to look like @sheg

OpenStudy (anonymous):

do one thing just in place of ? type 256

OpenStudy (anonymous):

sqrt{4}{256}

OpenStudy (anonymous):

\[\sqrt[4]{256}\]

OpenStudy (anonymous):

if i type that in the room will it show up the correct way/